# General Binding Affinity Guidance for Diffusion Models in Structure-Based Drug Design

### v2

@@ Abstract
keywords:
American Chemical Society,
[…]
{tocentry}
Synopsis Sentence: BADGER is a flexible guidance method that steers binding affinity toward desired values during molecular generation with diffusion models for structure-based drug design.
1.  Introduction
@@
Structure-based drug design (SBDD) is a fundamental task in drug discovery, aimed at designing ligand molecules that have a high binding affinity to the receptor protein pocket 1, 2. SBDD directly utilizes the three-dimensional structures of target proteins, enabling the design of molecules that can specifically interact with and influence the activity of these proteins, thus increasing the specificity and efficacy of potential drugs. A major challenge in SBDD is to design ligand molecules with desired properties, which includes properties like the binding affinity, as well as other physical and chemical properties. Conventional SBDD achieves this by conducting a “filtering” and “optimization” style workflow. During the filtering stage, a protein pocket is pre-selected and fixed, and a large database of ligand molecules is searched to find promising candidates with molecular docking. In the subsequent optimization stage, these candidate ligands are refined to improve desired properties using quantum mechanics or molecular mechanics (QM/MM) methods, combined with expert intuition and experience 3, 4, 5. However, this type of workflow faces several challenges. First, high-throughput experimental techniques or computational methods are both time-consuming and computationally demanding. Second, the search space for potential drug molecules is confined to the chemical database used in SBDD, limiting the diversity of candidates. Third, the optimization of candidate molecules post-identification is often influenced by human experience, which can introduce biases. These issues highlight the need for novel methods in SBDD to address these limitations effectively.
@@
Recent advances in machine learning, particularly in generative modeling, have provided a potentially computationally efficient alternative to the traditional SBDD approach. These developments can help overcome the limitations associated with the extensive ligand screening databases traditionally used in SBDD 6, 7, 8, 9, 10, 11, 12. Among the various types of generative models used for SBDD, diffusion models have shown some success in generating ligands that have high binding affinity to their target protein pockets 13, 6, 14, 15. In SBDD, diffusion model approaches model the continuous distribution of atom coordinates and discrete distribution of element types for both the ligand and protein. Schneuing et al. 15 and Guan et al. 6 first introduced diffusion models for SBDD by conditioning both the training and sampling stages of ligand molecule design on the protein pocket. Building on this idea, Guan et al. 14 developed a fragment-based strategy to improve ligand validity and binding affinity. Their approach decomposes ligands into fragments and initializes fragment positions using pre-designed priors before sampling. However, the effectiveness of this method strongly depends on the type and quality of the priors, which are tailored to specific pocket and ligand families, thereby limiting its generalizability to new systems. To further improve pocket–ligand binding affinity, Zhou et al. 16 proposed a filtering-based approach that incorporates physics-based predictors, such as AutoDock Vina’s scoring function (ADV energy function), during sampling. This method ranks and selects top candidates based on predicted binding affinity, but achieving substantial gains requires generating a large pool of ligands for filtering, which increases computational demands. Thus, developing a general and efficient strategy for integrating binding affinity into diffusion models for SBDD remains an open challenge.
@@
A key advantage of diffusion models is their ability to perform conditional sampling, where the model can be conditioned on specific properties during the training or sampling stages. This capability has been extensively validated in fields like image generation 13, 17. In the context of SBDD, applying conditional sampling strategies holds promise. By conditioning the model on desired ligand properties, diffusion models can facilitate the one-step design of ligands with optimal characteristics for a given protein pocket, offering a pocket agnostic, streamlined and highly targeted approach to drug design.
@@
There are two commonly used conditional sampling strategies in diffusion models: classifier guidance and classifier-free guidance, each applied at different stages of the diffusion process. Classifier guidance achieves conditional sampling by utilizing the gradients of separately trained classifiers to iteratively steer the sampled data toward regions with desired properties during the sampling stage. This approach is a plug-and-play post-training method that does not require retraining the diffusion model. In contrast, classifier-free guidance operates during the training phase. It involves training a single diffusion model that jointly models both conditional and unconditional distributions by randomly dropping the condition label during training. While this method requires modifications to the diffusion model’s training process, it eliminates the need for a separate classifier to provide guidance. Meanwhile, this method does not explicitly introduce additional computational cost during the sampling stage compared to unconditional diffusion sampling. In most image generation tasks, the condition for these methods is typically a class label represented by a multinomially distributed vector. However, in drug design, many relevant ligand properties, such as binding affinity, are continuous scalars distributed over specific intervals with physical meaning. Binding affinity is a critical measure of how effectively a ligand interacts with a protein pocket. In practice, binding affinity is often approximated using scoring functions like AutoDock Vina’s energy function (denoted as ADV energy function), which estimates binding free energy based on atomic interactions 4.
@@
[…]
@@ 1.2 Diffusion Models for Structure-based Drug Design
Recent advancements in generative modeling have been effectively applied to the SBDD task 23, 24, 25. The development of denoising diffusion probabilistic models 26, 17, 27, 13 has led to approaches in SBDD using diffusion models 6, 14, 16.
In the current literature of diffusion models for SBDD, both protein pockets and ligands are modeled as point clouds. In the sampling stage, protein pockets are treated as the fixed ground truth across all time steps, while ligands start as Gaussian noise and are progressively denoised. This process is analogous to image inpainting tasks, where protein pockets represent the existing parts of an “image,” and ligands are the “missing” parts that need to be filled in. Current approaches typically handle the ligand either as a whole entity 6, 15 or by decomposing ligands into fragments for sampling with pre-imposed priors 14, 16. In this work, we apply our guidance strategy to both of these methods.
The idea of diffusion-model-based SBDD is to learn a joint distribution between the protein pocket $P$ and the ligand molecule $M$. The spatial coordinates $x\in\mathbb{R}^{N\times 3}$ and atom features $v\in\mathbb{R}^{N\times K}$ are modeled separately by Gaussian $\mathcal{N}$ and categorical distributions $\mathcal{C}$, respectively, due to their continuous and discontinuous nature. Here $N$ is the number of atoms and $K$ is the number of element types. The forward diffusion process is defined as follows 6, 28:
@@ 1.3.1 Classifier Guidance
Classifier guidance 13 is a plug-and-play method that is straightforward to implement to fine-tune diffusion sampling. It involves decomposing a conditional distribution $P(x_{t}|y)$ into an unconditional distribution $P(x_{t})$ and a classifier term $P(y|x_{t})$ through Bayes’ Rule:
@@
To understand classifier guidance, consider that we are interested in maximizing the likelihood that the sampled $x_{0}$ belongs to class $y$. From a score-matching perspective 26, 17, the gradient of the log probability $P(x_{t}|y)$ with respect to $x_{t}$ is approximated and simplified through the following steps:
@@ 2.1.3 Guidance Strategy
$x_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}(x_{t}+\beta_{t}\nabla_{x_{t}}\log P(x_{t}))-\frac{\beta_{t}}{\sqrt{\alpha_{t}}}\omega\nabla_{\bm{x_{t}}}\mathcal{L}(\Delta G_{predict},\Delta G_{target})+\sqrt{\beta_{t}}z.$ (17)
@@
In diffusion models, the score term $\nabla_{x_{t}}\log P(x_{t})$ can either be parameterized in noise $\epsilon_{t}$ space or in $x_{0}$ space. We follow Guan et al. 6 and parameterize the score term in $x_{0}$ space, as this yields better performance when training an unconditional diffusion model for SBDD. The parametrization is shown in Eq. 18:
@@ 2.3 Dataset and Model Baselines
We test the methods on two benchmark datasets: CrossDocked2020 18 and PDBBind v2020 19. For CrossDocked2020, our data preprocessing and splitting procedures follow the same setting used in TargetDiff and DecompDiff 6, 14. Following Guan et al. 6, we filter 22.5 million docked protein-ligand complexes based on the criteria of low RMSD for the selected poses (< 1 Å) and sequence identity less than $30\%$. We select 100,000 complexes for training and 100 complexes for testing. For training the classifier used for guidance, both the previous training complexes and the test complexes are included for training. For evaluation, we sample 100 ligands from each pocket, resulting in a total of 10,000 ligands sampled for benchmarking.
To evaluate our method on higher-quality data, we also conduct experiments on PDBBind v2020 and CASF201629. Since CASF2016 is a subset of PDBBind, we use CASF2016 as the test set and the remaining portion of PDBBind v2020 for training. We clean the dataset by retaining only the entries whose elemental compositions are supported by AutoDock Vina, resulting in 12,126 training samples and 184 test samples.
@@
We benchmark the performance of our guidance methods on two state-of-the-art diffusion models for SBDD: TargetDiff 6 and DecompDiff 14. We benchmark the classifier guidance variant on TargetDiff and DecompDiff to show its plug-and-play feature. We then benchmark the classifier-free guidance variant on TargetDiff. For DecompDiff, we experiment with two types of priors used in their paper: the reference prior, which we denote as DecompDiff Ref, and the pocket prior, which we denote as DecompDiff Beta. We additionally include DiffSBDD 28, a recently proposed diffusion-based SBDD method that jointly models atomic coordinates and molecular topology. We include two other SBDD diffusion models as baselines: IPDiff 30, and BindDM 31. We also compare BADGER with DecompOpt 16, an optimization method built for diffusion models for SBDD. Specifically, for DecompOpt, we select the groups in Zhou et al. 16: TargetDiff + Optimization, which we denote as TargetDiff w/ Opt., and DecompDiff + Optimization, which we denote as DecompOpt. We also compare our results with non-diffusion SBDD models: liGAN 32, GraphBP 24, AR 25, Pocket2Mol 23.
3.  Results and Discussion
@@ 3.2 Binding Specificity Across Protein Pockets
Strong binding alone does not ensure useful molecular design; ligands should also bind selectively to their intended targets. Following Gao et al. 22, we identify a top-10 ligand for each pocket and compute its on-target affinity $\Delta G_{\text{on-target}}$. We then cross-dock the same ligand into five randomly selected pockets to obtain off-target affinities $\Delta G_{\text{off-target}}$. The resulting specificity score is:
@@ 3.3 Ligand–Protein Pose Quality
Binding affinity metrics do not capture geometric feasibility. Thus, we additionally evaluate steric clashes following Harris et al. 33. Steric clashes penalize overlapping van der Waals volumes and serve as an indicator of physically unrealistic poses.
@@ 4.  Conclusion
5.  Data and Software Availability
[…]
6.  Supporting Information
@@ 7.  Acknowledgements
S1.  Derivation for the classifier guidance variant conditional score expression
@@
S2.  Ablation on the type of energy function
@@
S3.  Algorithm for training the classifier
@@
S4.  Algorithm for classifier guidance sampling
@@
S5.  Algorithm for classifier-free guidance sampling
@@
S6.  Algorithm for training diffusion model for classifier-free guidance
@@
For the architecture of the noise network and the binding energy prediction classifier, we both adopt the Uni-Mol architecture following 34. We also perform an ablation using EGNN 35 as the classifier architecture. We outline the details for architecture in Tab. 5.
@@
The Equivariant Graph Neural Network (EGNN) is built based on Igashov et al. 36. The model contains two equivariant graph convolution layers. The total number of parameters for the model is $0.3$ million.
@@
The EGNN is trained using Adam 37, with learning rate = $5e^{-4}$, weight decay = 0, $\beta_{1}$ = 0.95, and $\beta_{2}$=0.999. We use the ReduceLROnPlateau scheduler with decaying factor = 0.5, patience = 2 and minimum learning rate = $1e^{-6}$. We use a Mean Squared Error (MSE) loss. We train the model for 20 epochs, and the loss drops to $0.1$. For the loss, we apply loss masking to get rid of the invalid data. Specifically, for any data with a ground truth binding affinity >0>0 kcal/mol, we set the loss for this data to be zero during training.
@@
The Transformer is built based on Zhou et al. 38. The model contains 10 attention layers. The total number of parameters for the model is $2.9$ million.
@@
The Transformer is trained by using Adam 37, with learning rate = $5e^{-4}$, weight decay = 0, $\beta_{1}$ = 0.95, and $\beta_{2}$=0.999. We use ReduceLROnPlateau scheduler with decaying factor = 0.5, patience = 2 and minimum learning rate = $1e^{-6}$. We use a Mean Squared Error (MSE) loss. We train the model for 20 epochs, and the loss drop down to $0.02$. For the loss, we apply loss masking to get rid of the invalid data. Specifically, for any data with a ground truth binding affinity >0>0 kcal/mol, we set the loss for this data to be zero during training.
@@
For classifier guidance, we use the pre-trained checkpoint of the diffusion model from Guan et al. 6 and Guan et al. 14 for TargetDiff and DecompDiff, respectively. We apply our guidance method on top of these trained models.
@@
We calculated QED, SA, and binding affinity using the same code base as in Guan et al. 6. Diversity is calculated as follows for the sampled ligands, following Guan et al. 6, Guan et al. 14:
@@
S8.  Other benchmark table
We provide benchmarks of our method with DecompOpt. Tab. 7 shows the benchmarking results with DecompOpt 16. According to Zhou et al. 16, DecompOpt and TargetDiff w/ Opt. sample 600 ligands for each pocket and select the top 20 candidates filtered by AutoDock Vina. To compare with these approaches, we sample 100 ligands for each pocket and select the top 20 candidates to compute the final binding affinity performance. The results show that BADGER outperforms DecompOpt by up to 50% in Vina Score, Vina Min, and Vina Dock.
Table 7: We benchmark binding affinity performance with DecompOpt 16 on the same test set with 100 pockets. To compare with DecompOpt and TargetDiff w/ Opt. under the same conditions, we sample 100 ligands for each pocket. We then select the top 20 candidates to compute the final binding affinity performance.
@@
DecompDiff Beta 14 -6.33 -7.56 -8.50 -8.88
Diff. + OPT. TargetDiff w/ Opt. 16 -7.87 -7.48 -7.82 -7.48
DecompOpt 16 -5.87 -6.81 -7.35 -7.72
@@
S9.  Extra results on binding affinity improvement
@@
S10.  Extra results on Redocking RMSD for sampled molecules
Redocking RMSD measures how closely the model-generated ligand pose matches the AutoDock Vina docked pose 33. A lower redocking RMSD suggests better agreement between the pose before and after redocking, indicating that BADGER more accurately mimics the docking score function. Fig. 6 compares redocking RMSD across models with and without BADGER. The results show that BADGER lowers the RMSD, improving the quality of the ligand poses sampled from diffusion model.
@@
S11.  Consistency between vina score and experimental measured binding energy
@@
S12.  Ablation of guidance strength $s$
@@
S13.  Benchmarking on Structural Validity, Specificity, Bond Angle, Bond Length and Chirality
We benchmark the molecule structural validity for all baselines as well as our methods, the results are shown in Tab. 9, to obtain the structural validity, we follow 28 and calculate the percentage of the sampled molecules that pass the rdkit sanitize molecule function. The results shows that adding BADGER on top of other diffusion baseline doesn’t affect the structural validity of the sampled molecules while also improved it’s binding affinity.
@@
We designed an evaluation procedure for specificity inspired by Gao et al. 22, with modifications to improve robustness. Specifically, for each of the 100 protein pockets in the test set, we selected the top 10 generated molecules based on Vina docking scores. These molecules were then cross-docked into 5 randomly selected off-target pockets (with a consistent seed for all models), resulting in a total of 5,000 cross-docking evaluations. We discarded any docking results with invalid (positive) scores, and computed the average difference in docking energy between the on-target and off-target settings. This Specificity Score quantifies specificity, with more negative values indicating greater difference hence greater selectivity for the intended target. As shown in Tab. 10, the classifier guidance framework significantly improves specificity, outperforming all baselines. We hypothesize that the addition of the classifier introduces richer structural conditioning signals, enabling the model to generate ligands that are more sensitive to the geometry and environment of the binding pocket. This observation is consistent with our earlier findings that guidance reduces steric clashes and improves redocking RMSD.
@@
S14.  Ablation on guidance with imperfect classifier
@@
S15.  Benchmarking the speed of guidance
@@
S16.  Analysis of Poor Binding in Outlier Pockets
@@
S17.  Potential Reasons for Improved Diffusion Guidance with Clean-Data Classifier Training

### v3

@@ Abstract
1 Introduction
@@
Structure-based drug design (SBDD) is a fundamental task in drug discovery, aimed at designing ligand molecules that have a high binding affinity to the receptor protein pocket [1, 2]. SBDD directly utilizes the three-dimensional structures of target proteins, enabling the design of molecules that can specifically interact with and influence the activity of these proteins, thus increasing the specificity and efficacy of potential drugs. A major challenge in SBDD is to design ligand molecules with desired properties, which includes properties like the binding affinity, as well as other physical and chemical properties. Conventional SBDD achieves this by conducting a “filtering” and “optimization” style workflow. During the filtering stage, a protein pocket is pre-selected and fixed, and a large database of ligand molecules is searched to find promising candidates with molecular docking. In the subsequent optimization stage, these candidate ligands are refined to improve desired properties using quantum mechanics or molecular mechanics (QM/MM) methods, combined with expert intuition and experience [3, 4, 5]. However, this type of workflow faces several challenges. First, high-throughput experimental techniques or computational methods are both time-consuming and computationally demanding. Second, the search space for potential drug molecules is confined to the chemical database used in SBDD, limiting the diversity of candidates. Third, the optimization of candidate molecules post-identification is often influenced by human experience, which can introduce biases. These issues highlight the need for novel methods in SBDD to address these limitations effectively.
@@
Recent advances in machine learning, particularly in generative modeling, have provided a potentially computationally efficient alternative to the traditional SBDD approach. These developments can help overcome the limitations associated with the extensive ligand screening databases traditionally used in SBDD [6, 7, 8, 9, 10, 11, 12]. Among the various types of generative models used for SBDD, diffusion models have shown some success in generating ligands that have high binding affinity to their target protein pockets [13, 6, 14, 15]. In SBDD, diffusion model approaches model the continuous distribution of atom coordinates and discrete distribution of element types for both the ligand and protein. Schneuing et al. [15] and Guan et al. [6] first introduced diffusion models for SBDD by conditioning both the training and sampling stages of ligand molecule design on the protein pocket. Building on this idea, Guan et al. [14] developed a fragment-based strategy to improve ligand validity and binding affinity. Their approach decomposes ligands into fragments and initializes fragment positions using pre-designed priors before sampling. However, the effectiveness of this method strongly depends on the type and quality of the priors, which are tailored to specific pocket and ligand families, thereby limiting its generalizability to new systems. To further improve pocket–ligand binding affinity, Zhou et al. [16] proposed a filtering-based approach that incorporates physics-based predictors, such as AutoDock Vina’s scoring function (ADV energy function), during sampling. This method ranks and selects top candidates based on predicted binding affinity, but achieving substantial gains requires generating a large pool of ligands for filtering, which increases computational demands. Thus, developing a general and efficient strategy for integrating binding affinity into diffusion models for SBDD remains an open challenge.
@@
A key advantage of diffusion models is their ability to perform conditional sampling, where the model can be conditioned on specific properties during the training or sampling stages. This capability has been extensively validated in fields like image generation [13, 17]. In the context of SBDD, applying conditional sampling strategies holds promise. By conditioning the model on desired ligand properties, diffusion models can facilitate the one-step design of ligands with optimal characteristics for a given protein pocket, offering a pocket agnostic, streamlined and highly targeted approach to drug design.
@@
There are two commonly used conditional sampling strategies in diffusion models: classifier guidance and classifier-free guidance, each applied at different stages of the diffusion process. Classifier guidance achieves conditional sampling by utilizing the gradients of separately trained classifiers to iteratively steer the sampled data toward regions with desired properties during the sampling stage. This approach is a plug-and-play post-training method that does not require retraining the diffusion model. In contrast, classifier-free guidance operates during the training phase. It involves training a single diffusion model that jointly models both conditional and unconditional distributions by randomly dropping the condition label during training. While this method requires modifications to the diffusion model’s training process, it eliminates the need for a separate classifier to provide guidance. Meanwhile, this method does not explicitly introduce additional computational cost during the sampling stage compared to unconditional diffusion sampling. In most image generation tasks, the condition for these methods is typically a class label represented by a multinomially distributed vector. However, in drug design, many relevant ligand properties, such as binding affinity, are continuous scalars distributed over specific intervals with physical meaning. Binding affinity is a critical measure of how effectively a ligand interacts with a protein pocket. In practice, binding affinity is often approximated using scoring functions like AutoDock Vina’s energy function (denoted as ADV energy function), which estimates binding free energy based on atomic interactions [4].
@@
We evaluate BADGER on the CrossDocked2020 [18] and PDBBind [19] datasets, demonstrating improvements in ligand binding affinity compared to baseline diffusion models. Beyond single-objective optimization, we extend BADGER to multi-property diffusion guidance, jointly optimizing for binding affinity, quantitative estimate of drug-likeness (QED) [20], and synthetic accessibility (SA) [21]. We further show that incorporating affinity guidance improves binding selectivity, as measured by the binding specificity score [22]. The code for this work will be publicly available at https://github.com/ASK-Berkeley/BADGER-SBDD.
@@ 1.2 Diffusion Models for Structure-based Drug Design
Recent advancements in generative modeling have been effectively applied to the SBDD task [23, 24, 25]. The development of denoising diffusion probabilistic models [26, 17, 27, 13] has led to approaches in SBDD using diffusion models [6, 14, 16].
In the current literature of diffusion models for SBDD, both protein pockets and ligands are modeled as point clouds. In the sampling stage, protein pockets are treated as the fixed ground truth across all time steps, while ligands start as Gaussian noise and are progressively denoised. This process is analogous to image inpainting tasks, where protein pockets represent the existing parts of an “image,” and ligands are the “missing” parts that need to be filled in. Current approaches typically handle the ligand either as a whole entity [6, 15] or by decomposing ligands into fragments for sampling with pre-imposed priors [14, 16]. In this work, we apply our guidance strategy to both of these methods.
The idea of diffusion-model-based SBDD is to learn a joint distribution between the protein pocket $P$ and the ligand molecule $M$. The spatial coordinates $x\in\mathbb{R}^{N\times 3}$ and atom features $v\in\mathbb{R}^{N\times K}$ are modeled separately by Gaussian $\mathcal{N}$ and categorical distributions $\mathcal{C}$, respectively, due to their continuous and discontinuous nature. Here $N$ is the number of atoms and $K$ is the number of element types. The forward diffusion process is defined as follows [6, 28]:
@@ 1.3.1 Classifier Guidance
Classifier guidance [13] is a plug-and-play method that is straightforward to implement to fine-tune diffusion sampling. It involves decomposing a conditional distribution $P(x_{t}|y)$ into an unconditional distribution $P(x_{t})$ and a classifier term $P(y|x_{t})$ through Bayes’ Rule:
@@
To understand classifier guidance, consider that we are interested in maximizing the likelihood that the sampled $x_{0}$ belongs to class $y$. From a score-matching perspective [26, 17], the gradient of the log probability $P(x_{t}|y)$ with respect to $x_{t}$ is approximated and simplified through the following steps:
@@ 2.1.3 Guidance Strategy
$x_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}(x_{t}+\beta_{t}\nabla_{x_{t}}\log P(x_{t}))-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\beta_{t}}{\sqrt{\alpha_{t}}}\omega\nabla_{\bm{x_{t}}}\mathcal{L}(\Delta G_{predict},\Delta G_{target})}+\sqrt{\beta_{t}}z.$ (17)
@@
In diffusion models, the score term $\nabla_{x_{t}}\log P(x_{t})$ can either be parameterized in noise $\epsilon_{t}$ space or in $x_{0}$ space. We follow Guan et al. [6] and parameterize the score term in $x_{0}$ space, as this yields better performance when training an unconditional diffusion model for SBDD. The parametrization is shown in Eq. 18:
@@ 2.3 Dataset and Model Baselines
We test the methods on two benchmark datasets: CrossDocked2020 [18] and PDBBind v2020 [19]. For CrossDocked2020, our data preprocessing and splitting procedures follow the same setting used in TargetDiff and DecompDiff [6, 14]. Following Guan et al. [6], we filter 22.5 million docked protein-ligand complexes based on the criteria of low RMSD for the selected poses (< 1 Å) and sequence identity less than $30\%$. We select 100,000 complexes for training and 100 complexes for testing. For training the classifier used for guidance, both the previous training complexes and the test complexes are included for training. For evaluation, we sample 100 ligands from each pocket, resulting in a total of 10,000 ligands sampled for benchmarking.
To evaluate our method on higher-quality data, we also conduct experiments on PDBBind v2020 and CASF2016[29]. Since CASF2016 is a subset of PDBBind, we use CASF2016 as the test set and the remaining portion of PDBBind v2020 for training. We clean the dataset by retaining only the entries whose elemental compositions are supported by AutoDock Vina, resulting in 12,126 training samples and 184 test samples.
@@
We benchmark the performance of our guidance methods on two state-of-the-art diffusion models for SBDD: TargetDiff [6] and DecompDiff [14]. We benchmark the classifier guidance variant on TargetDiff and DecompDiff to show its plug-and-play feature. We then benchmark the classifier-free guidance variant on TargetDiff. For DecompDiff, we experiment with two types of priors used in their paper: the reference prior, which we denote as DecompDiff Ref, and the pocket prior, which we denote as DecompDiff Beta. We additionally include DiffSBDD [28], a recently proposed diffusion-based SBDD method that jointly models atomic coordinates and molecular topology. We include two other SBDD diffusion models as baselines: IPDiff [30], and BindDM [31]. We also compare BADGER with DecompOpt [16], an optimization method built for diffusion models for SBDD. Specifically, for DecompOpt, we select the groups in Zhou et al. [16]: TargetDiff + Optimization, which we denote as TargetDiff w/ Opt., and DecompDiff + Optimization, which we denote as DecompOpt. We also compare our results with non-diffusion SBDD models: liGAN [32], GraphBP [24], AR [25], Pocket2Mol [23].
3 Results and Discussion
@@ 3.2 Binding Specificity Across Protein Pockets
Strong binding alone does not ensure useful molecular design; ligands should also bind selectively to their intended targets. Following Gao et al. [22], we identify a top-10 ligand for each pocket and compute its on-target affinity $\Delta G_{\text{on-target}}$. We then cross-dock the same ligand into five randomly selected pockets to obtain off-target affinities $\Delta G_{\text{off-target}}$. The resulting specificity score is:
@@ 3.3 Ligand–Protein Pose Quality
Binding affinity metrics do not capture geometric feasibility. Thus, we additionally evaluate steric clashes following Harris et al. [33]. Steric clashes penalize overlapping van der Waals volumes and serve as an indicator of physically unrealistic poses.
@@ 4 Conclusion
5 Data and Software Availability
The dataset and its processing procedure used in this work can be found at https://github.com/guanjq/targetdiff. Our source codes will be made publicly available at https://github.com/ASK-Berkeley/BADGER-SBDD.
6 Supporting Information
@@ 7 Acknowledgements
S1 Derivation for the classifier guidance variant conditional score expression
@@
S2 Ablation on the type of energy function
@@
S3 Algorithm for training the classifier
@@
S4 Algorithm for classifier guidance sampling
@@
S5 Algorithm for classifier-free guidance sampling
@@
S6 Algorithm for training diffusion model for classifier-free guidance
@@
For the architecture of the noise network and the binding energy prediction classifier, we both adopt the Uni-Mol architecture following [34]. We also perform an ablation using EGNN [35] as the classifier architecture. We outline the details for architecture in Tab. 5.
@@
The Equivariant Graph Neural Network (EGNN) is built based on Igashov et al. [36]. The model contains two equivariant graph convolution layers. The total number of parameters for the model is $0.3$ million.
@@
The EGNN is trained using Adam [37], with learning rate = $5e^{-4}$, weight decay = 0, $\beta_{1}$ = 0.95, and $\beta_{2}$=0.999. We use the ReduceLROnPlateau scheduler with decaying factor = 0.5, patience = 2 and minimum learning rate = $1e^{-6}$. We use a Mean Squared Error (MSE) loss. We train the model for 20 epochs, and the loss drops to $0.1$. For the loss, we apply loss masking to get rid of the invalid data. Specifically, for any data with a ground truth binding affinity >0>0 kcal/mol, we set the loss for this data to be zero during training.
@@
The Transformer is built based on Zhou et al. [38]. The model contains 10 attention layers. The total number of parameters for the model is $2.9$ million.
@@
The Transformer is trained by using Adam [37], with learning rate = $5e^{-4}$, weight decay = 0, $\beta_{1}$ = 0.95, and $\beta_{2}$=0.999. We use ReduceLROnPlateau scheduler with decaying factor = 0.5, patience = 2 and minimum learning rate = $1e^{-6}$. We use a Mean Squared Error (MSE) loss. We train the model for 20 epochs, and the loss drop down to $0.02$. For the loss, we apply loss masking to get rid of the invalid data. Specifically, for any data with a ground truth binding affinity >0>0 kcal/mol, we set the loss for this data to be zero during training.
@@
For classifier guidance, we use the pre-trained checkpoint of the diffusion model from Guan et al. [6] and Guan et al. [14] for TargetDiff and DecompDiff, respectively. We apply our guidance method on top of these trained models.
@@
We calculated QED, SA, and binding affinity using the same code base as in Guan et al. [6]. Diversity is calculated as follows for the sampled ligands, following Guan et al. [6, 14]:
@@
S8 Other benchmark table
We provide benchmarks of our method with DecompOpt. Tab. 7 shows the benchmarking results with DecompOpt [16]. According to Zhou et al. [16], DecompOpt and TargetDiff w/ Opt. sample 600 ligands for each pocket and select the top 20 candidates filtered by AutoDock Vina. To compare with these approaches, we sample 100 ligands for each pocket and select the top 20 candidates to compute the final binding affinity performance. The results show that BADGER outperforms DecompOpt by up to 50% in Vina Score, Vina Min, and Vina Dock.
Table 7: We benchmark binding affinity performance with DecompOpt [16] on the same test set with 100 pockets. To compare with DecompOpt and TargetDiff w/ Opt. under the same conditions, we sample 100 ligands for each pocket. We then select the top 20 candidates to compute the final binding affinity performance.
@@
DecompDiff Beta [14] -6.33 -7.56 -8.50 -8.88
Diff. + OPT. TargetDiff w/ Opt. [16] -7.87 -7.48 -7.82 -7.48
DecompOpt [16] -5.87 -6.81 -7.35 -7.72
@@
S9 Extra results on binding affinity improvement
@@
S10 Extra results on Redocking RMSD for sampled molecules
Redocking RMSD measures how closely the model-generated ligand pose matches the AutoDock Vina docked pose [33]. A lower redocking RMSD suggests better agreement between the pose before and after redocking, indicating that BADGER more accurately mimics the docking score function. Fig. 6 compares redocking RMSD across models with and without BADGER. The results show that BADGER lowers the RMSD, improving the quality of the ligand poses sampled from diffusion model.
@@
S11 Consistency between vina score and experimental measured binding energy
@@
S12 Ablation of guidance strength $s$
@@
S13 Benchmarking on Structural Validity, Specificity, Bond Angle, Bond Length and Chirality
We benchmark the molecule structural validity for all baselines as well as our methods, the results are shown in Tab. 9, to obtain the structural validity, we follow [28] and calculate the percentage of the sampled molecules that pass the rdkit sanitize molecule function. The results shows that adding BADGER on top of other diffusion baseline doesn’t affect the structural validity of the sampled molecules while also improved it’s binding affinity.
@@
We designed an evaluation procedure for specificity inspired by Gao et al. [22], with modifications to improve robustness. Specifically, for each of the 100 protein pockets in the test set, we selected the top 10 generated molecules based on Vina docking scores. These molecules were then cross-docked into 5 randomly selected off-target pockets (with a consistent seed for all models), resulting in a total of 5,000 cross-docking evaluations. We discarded any docking results with invalid (positive) scores, and computed the average difference in docking energy between the on-target and off-target settings. This Specificity Score quantifies specificity, with more negative values indicating greater difference hence greater selectivity for the intended target. As shown in Tab. 10, the classifier guidance framework significantly improves specificity, outperforming all baselines. We hypothesize that the addition of the classifier introduces richer structural conditioning signals, enabling the model to generate ligands that are more sensitive to the geometry and environment of the binding pocket. This observation is consistent with our earlier findings that guidance reduces steric clashes and improves redocking RMSD.
@@
S14 Ablation on guidance with imperfect classifier
@@
S15 Benchmarking the speed of guidance
@@
S16 Analysis of Poor Binding in Outlier Pockets
@@
S17 Potential Reasons for Improved Diffusion Guidance with Clean-Data Classifier Training